OpenStudy (anonymous):

The volumes of two spheres are 729 in3 and 27 in3. Find the ratio of their actual surface areas.

geerky42 (geerky42):

Volume of sphere: \(\Large\dfrac{4}{3}\pi r^3\) Surface area of sphere: \(\Large4\pi r^2\) Can you find radius for small and large sphere?